# Do Not (Always) Look Right: Investigating the Capabilities of Decoder-Based Large Language Models for Sequence Labeling

### 25 Jan 2024 v1

@@ 4 Experimental Setup
4.1 Datasets
We focus our experiments on IE tasks that can be framed as SL. Specifically, we choose three IE tasks: NER, ABSA with aspect term extraction and polarity subtasks, and a subtask of EE called trigger classification. Furthermore, since we tackle SL tasks, we include one more standard SL task unrelated to IE – text chunking. Dataset statistics are given in Table 1.
Dataset Train Validation Test
#Sent #Sent #Sent
CoNLL03 14041 3250 3453
ACE05 14672 873 711
ABSA Restaurants 2737 304 800
Table 1: Statistics for the datasets and their splits.
NER and Chunking.
For NER and text chunking, we choose CoNLL03, a standard and widely used benchmark (Tjong Kim Sang and De Meulder, 2003). We use a dataset version from “Hugging Face Datasets” (Lhoest et al., 2021) with an IOB2 sequence tagging scheme, which has train, validation, and test splits.
@@ 4.1 Datasets
We use the dataset from SemEval-2014 Task 4 and the “restaurants” domain (Pontiki et al., 2014) (ABSA Restaurants). We merge the first two ABSA subtasks, aspect term extraction (SB1) and aspect term polarity (SB2), into one SL task. We tokenized the original dataset with spaCy (Honnibal et al., 2020) and matched given character spans of aspect terms with token spans to create IOB2 tags for sentences (13 aspect terms could not be matched automatically in the whole dataset, so we ignored these aspect terms). Training and test sets were given. Following prior work (Wang et al., 2021), we randomly sample $10\%$ of the sentences for a validation set.
@@
The ACE05 dataset (Doddington et al., 2004) is a widely used trigger classification dataset. Trigger classification encapsulates two SL tasks: trigger identification, i.e., finding spans of trigger tokens and classifying those spans. We use the English train, validation, and test splits obtained with the standard ACE pre-processing tool. [^1] https://bit.ly/ace2005-preprocessing We use this tool to obtain sentences and tokens and create IOB2 tags.
@@
We evaluate all tasks with micro F1 score on IOB2 tag predictions with strict matching using “seqeval” (Nakayama, 2018), where the predicted output span must exactly match the expected output span. We evaluate only the predictions on the first token of the tokenized words from the input sequence to obtain the same number of predictions as there are target IOB2 labels.
@@
A special pre-processing of each SL dataset is required to train LLMs for IT. We create instructions similarly to the ones used for NER by Wang et al. (2022b), although we require a more strict output response from decoder-based LLMs, similar to the output format used by Wang et al. (2023b). Instruction templates are created by parsing IOB2 tags to create desired outputs. Examples of instruction tuning examples created with those templates are given in Appendix A.1. To ensure fair evaluation consistent with models trained for SL, we map response spans of instruction-tuned models to IOB2 tags with various heuristics. We employ greedy span-based matching of predicted spans and their types with input tokens. If no predictions are made, all predicted spans do not align with input tokens, or if an error arises during matching due to output generation stochasticity, we treat all these scenarios as if the “O” tag was predicted for every input token. This approach is similar to the evaluation of Wang et al. (2022b).
@@
Figure 1: Micro F1 SL scores with decoder-based LLMs and different unmasking configurations (sorted by Gray code starting with all decoder layers masked – code $0$). Upper row plots show Llama2-7B model results, while lower row plots show Mistral-7B model results on validation and test sets of four SL tasks (left to right, one dataset per task). All results are averages of five runs with different seeds. The shaded area represents the standard deviation.
@@ 4.2 Training Details and Hyperparameters
To fight out-of-memory issues, we pre-process all SL datasets to have a maximum tokenized sequence length of $128$. This, however, poses no problem because there are less than $10$ sentences for each dataset and split that end up truncated independent of the tokenizer used. We use the end-of-sequence token for the models with no padding token to pad the sequences to the longest example in the batch. We randomly sample examples for training depending on the seed. For MLM-based encoder experiments, we use strong RoBERTa-base and RoBERTa-large PLMs with 125 and 355 million parameters, respectively (Liu et al., 2019). We choose “open” LLMs for unmasking experiments and IT: Llama2-7B (Touvron et al., 2023) and Mistral-7B (Jiang et al., 2023), both with seven billion parameters. Since RoBERTa requires tokenization with added prefix space, we enforce the same for “open” LLMs’ tokenizers. For SL experiments, we employ a classic softmax SL head on top of the pre-trained models. Models and their implementations are taken from “Hugging Face Transformers” Wolf et al. (2020). If the SL head is not implemented, we add our implementation as it was implemented for RoBERTa models.
@@
We use quantized LoRA (QLoRA) (Hu et al., 2021; Dettmers et al., 2023) for fine-tuning all models to ensure a fair comparison between the smallest and the largest models and make fine-tuning possible under constrained computing requirements. LoRA is applied to query and value attention matrices inside each encoder or decoder block with a fixed rank of $r=64$, a scaling parameter of $\alpha=16$, and a dropout probability of $p=0.1$. This way, only decomposed query and value matrices, and the SL head parameters are optimized with cross-entropy loss, yielding a drastic trainable parameter reduction per model. The cross-entropy loss is adjusted to consider only the first token of each tokenized word from the input sequence. The models are trained in bfloat16 precision, with loaded pre-trained weights in 4-bit NormalFloat data type, and we use double quantization. With QLoRA and the mentioned quantization configuration, all models can fit into 40GB of GPU memory of “Ampere A100” and are trained with a consistent batch size of $16$ per experiment. We train the models with paged 8-bit AdamW (Loshchilov and Hutter, 2017) optimizer to handle the memory spikes Dettmers et al. (2023). The parameters of AdamW are fixed to $\beta_{1}=0.9,\beta_{2}=0.95,\epsilon=1\mathrm{e}{-5},\lambda=0.1$. For the learning rate scheduler, we choose cosine annealing scheduler (Loshchilov and Hutter, 2016). We apply gradient clipping to $1.0$, gradient accumulation with four steps, and gradient checkpointing. We use a consistent learning rate of $2\mathrm{e}{-4}$ across all experiments and fine-tune PLMs or LLMs over a fixed number of five epochs. All results are averages of five runs with different seeds ($seeds=[120,121,122,123,124]$), and we always pick the last model for each seed. We do not tune the hyperparameters due to the computing constraints but fix them to the mentioned recommended values in advance.
@@
We use the PLMs and LLMs with CLM heads from “Hugging Face Transformers” for IT experiments. Training is also performed with QLoRA based on cross-entropy loss. All hyperparameters are identical to SL experiments except for the batch size. The batch size depends on the maximum sequence length for packing the dataset examples. Short examples are packed together in the same input to increase training efficiency. This length is set to $512$ for the ABSA Restaurants dataset and to $1024$ for other datasets since ABSA Restaurants has shorter instruction examples on average compared to other datasets. It also has the lowest number of examples in the training set (cf. Table 1). We observe that ABSA models need double the epochs ($10$ epochs) to learn to extract aspect terms and their polarity, and we hypothesize that this is due to the low number of training examples compared to other datasets which successfully solve the tasks with five epochs of supervised instruction fine-tuning. We generate the outputs for evaluation with default generation settings for the Llama2-7B model. We only decrease the total maximum length of the input instruction prompt with newly generated tokens to $1024$ to speed up the generation. The same generation config is used for Mistral-7B.
@@ 5 Results
Model CoNLL03 NER ABSA Restaurants ACE05 Trigger Clf. CoNLL03 Chunking Llama2-7B-SL Valid F1 Test F1 $\rho$ Valid F1 Test F1 $\rho$ Valid F1 Test F1 $\rho$ Valid F1 Test F1 $\rho$ Unmask Config. $0$ 0.788 0.744 0.999 0.430 0.482 0.994 0.591 0.574 0.806 0.875 0.864 0.998 $1$ 0.936 0.908 0.345 0.402 0.635 0.584 0.892 0.880 Best 0100 - - 0.488 0.546 - - - - 0101 - - - - - - - 0.885 - 0110 0.947 0.920 - - - - 0.898 - 0111 - - - - 0.653 - - - 1100 - - - - - 0.616 - - Mistral-7B-SL Valid F1 Test F1 $\rho$ Valid F1 Test F1 $\rho$ Valid F1 Test F1 $\rho$ Valid F1 Test F1 $\rho$ Unmask Config. $0$ 0.789 0.756 0.999 0.484 0.553 0.983 0.607 0.584 0.899 0.882 0.872 0.994 $1$ 0.953 0.927 0.581 0.641 0.668 0.626 0.904 0.892 Best 0110 - - 0.597 0.659 - - - - 0111 0.956 - - - - - - - 1101 - - - - - 0.639 - - 1110 - 0.927 - - 0.676 - - - RoBERTa-base-SL 0.897 0.883 - 0.313 0.369 - 0.609 0.508 - 0.889 0.877 - RoBERTa-large-SL 0.924 0.900 - 0.403 0.474 - 0.698 0.628 - 0.891 0.877 - Llama2-7B-IT 0.778 0.771 - 0.523 0.608 - 0.375 0.347 - 0.833 0.818 - Mistral-7B-IT 0.897 0.887 - 0.646 0.733 - 0.477 0.461 - 0.873 0.860 -
Table 2: Validation and test micro F1 SL scores for Llama2-7B and Mistral-7B models with different unmasking configurations ($0$, $1$, and other configurations which surpass the F1 score of $1$ configuration over four SL datasets – denoted as “Best”) are in the upper part of the table. The results for SL RoBERTa baselines and instruction tuning (IT) baselines are given in the lower part of the table. The best results by dataset and evaluation set are in bold. For Llama2-7B and Mistral-7B, we report the Pearson correlation coefficient between validation and test unmasking configurations ($\rho$). All results are averages of five runs with different seeds.
@@ 5.1 Layer Group Unmasking Results
Figure 1 shows the validation and test set performance of Llama2-7B and Mistral-7B for different unmasking configurations and four SL tasks. We observe some interesting phenomena. (1) In most cases, removing the CM in as little as one layer group significantly boosts the micro F1 SL score compared to keeping the CM in all layers (code $0000$) during fine-tuning. (2) It is rarely the case that removing CM from all layer groups (code $1111$), compared to removing it only from some layer groups, yields the highest score on validation and test set. (3) Depending on the task, the boosts from the CM removal can vary (highest boosts for NER, lowest for text chunking) and deviate from minimal (NER) to substantial (trigger classification). (4) Mistral-7B shows the superiority over Llama2-7B model across tasks and best configurations. (5) Performance on validation and test sets highly correlates across unmasking configurations and datasets for a fixed model.
Unmasking gains are the highest for the NER task, and the results are generally the cleanest for CoNLL03 data (NER and text chunking tasks). Refining the layer-wise CM removal inside individual layers could result with even larger micro F1 score boosts. The performance of Llama2-7B on the ABSA Restaurants dataset shows a peculiar trend where better performance is achieved when CM is kept in all layer groups compared to CM removal. It is as if allowing the model to see a broader context hurts the performance. This trend is not present in any other dataset and model combination. Furthermore, the ABSA Restaurants dataset is the only one with higher scores on the test set than on the validation set, which could be attributed to taking a random sample of training data as validation set. The deviations of results across evaluation sets are the biggest on ABSA and trigger classification tasks which could be due to the overall low number of examples per evaluation set for those tasks ($<1000$).
@@ 5.2 Comparison With Baselines
To compare the results between different unmasking configurations and strong encoder or instruction-tuned baselines, we present the F1 scores in Table 2. The results are compared against MLM-based RoBERTa PLMs fine-tuned for SL task and instruction-tuned Llama2-7B and Mistral-7B LLMs. We compare the standard $0000$ and $1111$ unmasking configurations with the best configurations overall per model and task. We also report the Pearson correlation coefficient $\rho$ between all validation and test unmasking configurations. For all SL tasks and the Llama2-7B model, we observe consistent improvement over the $1111$ unmasking configuration. For Mistral-7B, similar holds, except for the chunking tasks where $1111$ unmasking configurations yield the highest F1 scores. IT is most beneficial for the ABSA task, which aligns with findings from Scaria et al. (2023) and outperforms any unlocking configuration on validation and test sets. RoBERTa-large baseline surpassed the performance of all other models on the ACE05 validation set but failed to do so on the test set. RoBERTa baselines generally achieve high scores on SL tasks, except for ABSA. Trainin with QLoRA in combination with strategic layer unmasking creates high-performing and compact models, requiring a small number of additional parameters trained for each SL task. The ACE05 dataset exhibits the lowest correlation between performance on the validation and test sets. Nevertheless, the generally high correlation across all datasets indicates that the validation set can effectively be utilized to select the optimal unmasking configuration for each model and task.
@@ 5.3 Comparison With SOTA IE Models
Our results are competitive with the SOTA approaches. However, it is challenging to ensure a fair comparison due to differences in training and evaluation. For example, the SOTA result for ConLL03 NER is the F1 score of $94.6$, reported by Wang et al. (2021). However, this result is obtained using a model trained on the merged training and validation set. Moreover, important details of the evaluation setup were not specified (micro vs. macro F1 score, token- vs. span-based evaluation, and evaluating the model prediction on the first token of tokenized words from the input sequence vs. on all tokens). Furthermore, the authors whose work we build upon, Li et al. (2023), report SOTA results on CoNLL03 NER by removing the CM from LLama2. The score they report is $93.19$, which is competitive with Wang et al. (2021), but not SOTA. Upon code inspection, we found out that they do not use QLoRA, so they cut off sentences longer than $64$ tokens to fit the data and model into GPU memory. In contrast, we managed to keep the sequence length at $128$ tokens and cut only a few sentences per task ($<10$). These decisions affect the overall performance. Scaria et al. (2023) achieve SOTA on ABSA Restaurants dataset, reaching F1 score of $92.76$ on aspect term extraction with IT. However, they do not report the results on combined aspect term extraction and polarity but separately treat the two tasks. The SOTA F1 score results of $69.8$ on ACE05 for trigger classification are achieved with the DeepStruct model Wang et al. (2022a).
5.4 Pre-training Small Encoders and Decoders

### 27 Feb 2024 v2

@@ 4 Experimental Setup
4.1 Sequence Labeling Tasks
We focus our experiments on IE tasks framed as SL: NER, ABSA with aspect term extraction and polarity subtasks, and the trigger classification (TC) subtask of EE. We also include text chunking (shallow parsing), which, although not an IE task, is considered a prototypical SL task. Dataset statistics are shown in Table 1. For each dataset, we calculate its right-side dependency relations ratio (RDRR), defined as the ratio of right-side dependency relationships to the total number of left-side and right-side relationships counted for all labeled spans in the training set. Essentially, this metric indicates the degree to which labeled spans depend on the context to their right. We use spaCy (Honnibal et al., 2020) dependency parser to obtain the dependency relations.
Dataset Train Validation Test Total RDRR CoNLL03 NER 14041 3250 3453 20744 0.593 CoNLL03 Chunking 0.518 ACE05 14672 873 711 16256 0.411 Rest14 2737 304 800 3841 0.397
Table 1: Statistics for the datasets and their splits. We show number of sentences per split, total number of sentences, and right-side dependency relations ratio.
NER and Text Chunking.
For NER and text chunking, we choose CoNLL03, a standard and widely used benchmark (Tjong Kim Sang and De Meulder, 2003). We use the version from Hugging Face Datasets (Lhoest et al., 2021) with an IOB2 sequence tagging scheme, which has a predefined train, validation, and test split.
@@ 4.1 Sequence Labeling Tasks
We use the data from SemEval-2014 Task 4 and the restaurants domain (Pontiki et al., 2014) (Rest14). We merge the first two ABSA subtasks, aspect term extraction (ATE) and aspect term polarity (ATP), into one SL task (ATE+ATP). We tokenize the dataset with spaCy (Honnibal et al., 2020) and match given character spans of aspect terms with token spans to obtain IOB2 tags (we discard 13 aspect terms that could not be matched in this way). Training and test split were predefined. Following prior work (Wang et al., 2021), we randomly sample 10% of the sentences for the validation set.
@@
The ACE05 dataset (Doddington et al., 2004) is a widely used TC dataset. The TC task combines two EE tasks into a single SL task: trigger identification, i.e., finding spans of tokens constituting the event predicate and classifying them. We use the English train, validation, and test split obtained with the standard ACE pre-processing tool. [^2] https://bit.ly/ace2005-preprocessing We use this tool to obtain sentences and tokens and create IOB2 tags.
@@
We evaluate all tasks with micro F1 score on IOB2 tag predictions with strict matching using seqeval (Nakayama, 2018), where the predicted output span must exactly match the expected output span. We evaluate only the predictions on the first token of the tokenized words from the input sequence to obtain the same number of predictions as there are target IOB2 labels.
@@
To train LLMs using IT, each dataset needs to be further pre-processed to obtain instruction prompts. We form instructions similar to the ones used for NER by Wang et al. (2022b), although we require a more strict output response from decoder-only LLMs, in line with the output format used by Wang et al. (2023b). We create instruction prompts by parsing IOB2 tags to create desired outputs. See Appendix A.3 for instruction tuning examples. To ensure a fair evaluation consistent with models fine-tuned directly for SL, we heuristically map response spans of instruction-tuned models to IOB2 tags. We employ greedy span-based matching of predicted spans and their types with input tokens, similar to Wang et al. (2022b). We treat all cases in which no predictions are made, or all predicted spans do not align with input tokens, or an exception arises during matching due to output generation stochasticity, as if the O tag was predicted for every input token.
@@
Figure 1: Micro F1 SL scores with decoder-based LLMs and different unmasking configurations (sorted by Gray code starting with all decoder layers masked – configuration $0$). Upper row plots show Llama2-7B model results, while lower row plots show Mistral-7B model results on validation and test sets of four SL tasks (left to right, one dataset per task). All results are averages of five runs. The shaded area corresponds to standard deviation.
@@ 4.2 Training Details and Hyperparameters
Models.
We choose open LLMs for unmasking experiments and IT: Llama2-7B (Touvron et al., 2023) and Mistral-7B (Jiang et al., 2023), both with 7B parameters. For encoders, we use RoBERTa-base and RoBERTa-large PLMs with 125 and 355 million parameters, respectively (Liu et al., 2019). For SL experiments, we employ a classic softmax SL head on top of the pre-trained models. Model implementations and weights are taken from Hugging Face Transformers Wolf et al. (2020). If the SL head is not implemented, we add our implementation mimicking the one for RoBERTa models. See Appendix A.1 for more details.
Optimization and Pre-processing.
@@
We use quantized LoRA (QLoRA) (Hu et al., 2021; Dettmers et al., 2023) for fine-tuning all models to ensure a fair comparison between the smallest and the largest models and enable fine-tuning under constrained computing resources. QLoRA is applied to query and value attention matrices inside each encoder or decoder block with a fixed rank of $r=64$, a scaling parameter of $\alpha=16$, and a dropout probability of $p=0.1$. This way, only decomposed query and value matrices, along with the SL head parameters, are optimized with cross-entropy loss, yielding a drastic trainable parameter reduction per model. The models are trained in bfloat16 precision, with loaded pre-trained weights in 4-bit NormalFloat data type, and we use double quantization. Using this setup, we were able to fit all models into 40GB of GPU memory of Ampere A100 and are trained with a consistent batch size of 16 per experiment. To use this batch size across models, we pre-process all datasets to a maximum tokenized sequence length of 128. This cutoff is optimal as there are <10 sentences for each dataset and split that end up truncated independent of the tokenizer used. We pad the sequences to the longest example in the batch and randomly sample examples for training depending on the seed. We train the models with paged 8-bit AdamW (Loshchilov and Hutter, 2017) optimizer to handle the memory spikes Dettmers et al. (2023). The parameters of AdamW are fixed to $\beta_{1}=0.9,\beta_{2}=0.95,\epsilon=1\mathrm{e}{-5},\lambda=0.1$. For the learning rate scheduler, we choose cosine annealing scheduler (Loshchilov and Hutter, 2016). We apply gradient clipping set to $1.0$, gradient accumulation with four steps, and gradient checkpointing. We use a consistent learning rate of $2\mathrm{e}{-4}$ across all experiments and fine-tune models over a fixed number of five epochs. All results are averages of five runs with different seeds, and we always pick the last model for each seed.
Instruction Tuning Details.
@@
We use Llama2-7B and Mistral-7B with CLM heads for IT. Training is done with QLoRA based on cross-entropy loss. All hyperparameters are inherited from SL experiments except for the batch size. The batch size depends on the maximum sequence length for packing the dataset examples. Short examples are packed together in the same input to increase training efficiency. This length is set to 512 for the Rest14 dataset and 1024 for other datasets since Rest14 has shorter instruction examples on average and has the smallest training set (cf. Table 1). We observe that ATE+ATP models need double the epochs (10 epochs) to learn the task. We hypothesize this is due to the low number of training examples compared to other datasets.
@@ 5.1 Layer Group Unmasking Results
Figure 1 shows the validation and test evaluation sets performance of Llama2-7B and Mistral-7B for different unmasking configurations and four SL tasks. We observe a number of interesting phenomena: (1) In most cases, removing the CM in as little as one layer group significantly boosts the F1 score compared to keeping the CM in all layers (configuration $0$); (2) It is rarely the case that removing CM from all layer groups (configuration $1$), compared to removing it only from some layer groups, yields the highest score on evaluation sets; (3) Depending on the task, the boosts from the CM removal can vary (highest for NER, lowest for chunking) and deviate from minimal (NER) to substantial (TC); (4) Mistral-7B shows superior performance over Llama2-7B model across tasks and best configurations; (5) The shapes of validation and test curves show high overlap across configurations for a fixed model and dataset.
[…]
Model CoNLL03 NER Rest14 ATE+ATP ACE05 Trigger Clf. CoNLL03 Chunking Llama2-7B-SL Valid F1 Test F1 Valid F1 Test F1 Valid F1 Test F1 Valid F1 Test F1 $\rho=0.999$ $\rho=0.994$ $\rho=0.806$ $\rho=0.998$ Unmask Config. $0$ 0.788 0.744 0.430 0.482 0.591 0.574 0.875 0.864 $1$ 0.936 0.908 0.345 0.402 0.635 0.584 0.892 0.880 Best 0100 0.488 0.546 0101 0.885 0110 0.947 0.920 0.898 0111 0.653 1100 0.616 Mistral-7B-SL Valid F1 Test F1 Valid F1 Test F1 Valid F1 Test F1 Valid F1 Test F1 $\rho=0.999$ $\rho=0.983$ $\rho=0.899$ $\rho=0.994$ Unmask Config. $0$ 0.789 0.756 0.484 0.553 0.607 0.584 0.882 0.872 $1$ 0.953 0.927 0.581 0.641 0.668 0.626 0.904 0.892 Best 0110 0.597 0.659 0111 0.956 1101 0.639 1110 0.927 0.676 RoBERTa-base-SL 0.897 0.883 0.313 0.369 0.609 0.508 0.889 0.877 RoBERTa-large-SL 0.924 0.900 0.403 0.474 0.698 0.628 0.891 0.877 Llama2-7B-IT 0.778 0.771 0.523 0.608 0.375 0.347 0.833 0.818 Mistral-7B-IT 0.897 0.887 0.646 0.733 0.477 0.461 0.873 0.860
Table 2: Validation and test micro F1 SL scores for Llama2-7B and Mistral-7B models with various unmasking configurations ($0$, $1$, and other configurations which surpass the F1 score of $1$ over SL datasets – denoted as Best) are in the upper table part. The results for SL encoders and IT baselines are in the lower table part. The best results by dataset and evaluation set are in bold. For Llama2-7B and Mistral-7B, we report the Pearson correlation coefficient between validation and test unmasking configurations ($\rho$). All results are averages over five runs.
@@ 5.2 Comparison with Baselines
Table 2 shows the performance scores of the best unmasking configurations and the strong encoder and IT baselines. We compare against RoBERTa PLMs fine-tuned for SL task and instruction-tuned Llama2-7B and Mistral-7B LLMs. The standard $0$ and $1$ configurations are compared with the best configurations per model and task. We also report the Pearson correlation coefficient $\rho$ between all validation and test unmasking configurations. For all SL tasks and the Llama2-7B model, we observe a consistent improvement of best configurations over configuration $1$. Similar holds for Mistral-7B, except for chunking, where $1$ yields the highest F1 scores. Conforming to the findings from Scaria et al. (2023), IT is most beneficial for the ATE+ATP task, outperforming any unmasking configuration on evaluation sets. RoBERTa-large surpasses all other models on the ACE05 validation set but fails to do so on the test set. RoBERTa baselines achieve high scores on SL tasks, except for ATE+ATP. Training with QLoRA combined with layer group unmasking creates high-performing and compact models, requiring a small number of additional parameters trained for each SL task. Correlations between configurations on evaluation sets are high for each task, while models trained on ACE05 exhibit the lowest $\rho$. A high overall $\rho$ indicates that the optimal unmasking configuration can be determined using the validation set.
@@ 5.3 Comparison with SOTA IE Models
Our results are competitive with SOTA. However, a fair comparison is challenging due to differences in training and evaluation. For example, the SOTA F1 score for CoNLL03 NER, reported by Wang et al. (2021), is $0.946$. This result, however, was obtained using a model trained on the merged training and validation sets. Moreover, important details of the evaluation setup were not specified (micro vs. macro F1 score, token- vs. span-based evaluation, and evaluating the model prediction on the first token of tokenized words from the input sequence vs. on all tokens). Further, the authors whose work we build upon, Li et al. (2023), report SOTA results on CoNLL03 NER. The score they report is $0.932$, which is competitive, but not SOTA. Upon code inspection, we found that they truncate sequences longer than 64 tokens to fit the data into GPU memory. In contrast, we managed to keep the sequence length at 128 tokens. These decisions have a significant impact on overall performance. Scaria et al. (2023) achieve SOTA F1 score of $0.928$ on Rest14 dataset and ATE task with IT. Yang and Li (2021) report SOTA macro F1 of $0.863$ on ATP task with DeBERTa model He et al. (2020). The SOTA F1 score of $69.8$ on ACE05 for TC is achieved by Wang et al. (2022a).
5.4 Investigating the Effect of Scale